OpenStudy (anonymous):

Please help me find the function rule for this table!!! Please explain if you can I will love you forever! Thank you! xx x - 2, 4, 6 y - 1, 0, -1

jimthompson5910 (jim_thompson5910):

as x goes up by 2, y decreases by 1 so this suggests we have a linear function with a slope of m = rise/run = -1/2

jimthompson5910 (jim_thompson5910):

now use any point in the table to find the equation let's say we use (4,0) y = mx + b y = (-1/2)x + b .... plug in the slope 0 = (-1/2)(4) + b ... plug in the point we'll use 0 = -2 + b keep going to solve for b and tell me what you get

OpenStudy (anonymous):

Gah Im confused

jimthompson5910 (jim_thompson5910):

0 = -2 + b is the same as 0 = b - 2 so how would you move that -2 over

jimthompson5910 (jim_thompson5910):

or you think of it as SOME NUMBER - 2 = 0 what goes in the place of "SOME NUMBER" ?

OpenStudy (anonymous):

I understand. Lol well thanks a lot. xx

jimthompson5910 (jim_thompson5910):

so what did you get for b